Find the product, using suitable properties:
 (–57) × (–19) + 57

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given equation is in the form of Distributive law of Multiplication over Addition.
= a × (b + c) = (a × b) + (a × c)
= (57 × 19) + (57 × 1)
= 57 [19 + 1]
= 57 × 20
= 1140

Thanks